Suppose that each week, you deposit $44 into a savings account whose annual rate is 2.8% with weekly compounding. How much will you have in the account after 9 years?
Answer #1

Weekly payment=$44
Annual rate=2.8% so weekly rate=2.8%/52=0.000538462

Total number of periods for weekly compounding is (9 years)*(52 weeks in a year)=9*52=468
We can determine the future value using excel,
As the payments are cash outflows, we have taken it as negative


So, the future value is $23,411.86